OpenStudy (anonymous):

$8,000 is invested in two different accounts earning 3% and 4% interest. If the two accounts earn a total of $295 in interest, how much money was invested at 3%? A $2,500 B $3,500 C $5,500

OpenStudy (anonymous):

@amandaemiko Let A be the amount invested at 3%, B the amount at 4%. Total value: A+B=8000 Interest: Multiply eq. 1 by -3 and add to eq. 2, - 3A - 3B + 3A + 4B = -24000 + 29500 B=5500 Then from eq. 1, A + 5500 = 8000 A = 2500 so A is the answer